OpenStudy (anonymous):

Medal Solve. (Round to the nearest hundredth, if necessary.) A bank requires a monthly payment of $24.09 on a $2100 loan. At the same rate, find the monthly payment on a $52,500 loan. $ per month

OpenStudy (anonymous):

So (24.09/2100) is the ratio of payment required per dollar loan. So we have \[(24.09/2100) = (x/52500)\] Solve for x. Done.

OpenStudy (anonymous):

0.9636

OpenStudy (anonymous):

m i right?

OpenStudy (anonymous):

No

OpenStudy (calculusxy):

\[\frac{ 24.09 }{ 2100 } = \frac{ x }{ 52,500 }\]

OpenStudy (calculusxy):

You should thus have: \[52,500(24.09)=2100x\]

OpenStudy (anonymous):

602.25

OpenStudy (calculusxy):

Yes

OpenStudy (anonymous):

Yes, that's correct

OpenStudy (anonymous):

is that rounded to nearest hundredth?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

ok thank you everyone :)

OpenStudy (calculusxy):

You're welcome :)